OpenStudy (hellokitty17):

Choose the equation below whose axis of symmetry is x = 2. y = x2 + 4x + 2 y = x2 − 4 y = x2 − 2 y = x2 − 4x + 2

OpenStudy (hellokitty17):

EXPLAIN THE ANSWER!!

oregonduck (oregonduck):

x2 - 4x + 2 cuzz f(x-4)= (x-4)^2 expand and (x-4)(x-4) = that :P